OpenStudy (anonymous):

FREE MEDALS!!! if you answer! Tom and Tim are standing together in the park. Tom is 6 feet tall and his shadow is 4 feet long. How tall is Tim if his shadow is 2 feet long ? 3 feet 4 feet 6 feet 8 feet

OpenStudy (amistre64):

6 is to 4 as x is to 2 x = 3

OpenStudy (saifoo.khan):

3fts.

OpenStudy (amistre64):

tim must be a dog of some sorts

myininaya (myininaya):

3

OpenStudy (saifoo.khan):

lol @amistre.

OpenStudy (anonymous):

3feet

OpenStudy (anonymous):

6/4 x/2 Cross multiply. 6 x 2 = 12 4x = 4x 4x = 12 12 divided by 4 = 3 3 feet.

OpenStudy (anonymous):

hahahaha

OpenStudy (anonymous):

free medals?

OpenStudy (bahrom7893):

he's not a dog.. he's a midget..

OpenStudy (anonymous):

Or a little kid.

OpenStudy (amistre64):

now now; we all know the term midget is politically incorrect ;)

OpenStudy (anonymous):

2 Tan[ArcTan[6/4]] = 3

OpenStudy (anonymous):

....Rob put it in simple form.

OpenStudy (anonymous):

\[2 \text{Tan}[\text{ArcTan}[6/4]]=2*\frac{3}{2}=3 \]